### v2

# Efficient Reasoning

Abstract
[…]
1 Introduction
Figure 1: Test-time scaling; the horizontal axis represents relative wall clock duration on an AMD MI300X GPU averaged across 553 GenEval Ghosh et al. [2023] prompts and the vertical axis represents the Overall GenEval score. VQAScore Lin et al. [2024] is used as verifier and FLUX-Schnell Labs et al. [2025] as generator for all methods. The Verifier-Threshold method is consistently more efficient and delivers higher scores compared to the baseline FlowRBF Kim et al. [2025] as well as other methods.
Image generation using AI has become a popular application with several models DeepMind [2025], OpenAI [2025], Apple Machine Learning Research [2025], Midjourney [2025], Stability AI [2024] being widely used. However, even state-of-the-art models often struggle to accurately follow detailed prompt instructions Ghosh et al. [2023], Huang et al. [2023]. To address this limitation, the vision community has drawn inspiration from the language domain by incorporating test-time compute Snell et al. [2024] and reasoning Guo et al. [2025], leading to the development of analogous approaches for vision tasks Ma et al. [2025], Zhuo et al. [2025]. However, reasoning-based methods inherently demand computational power and time during inference, which becomes costly at large scales Jin et al. [2025] or on resource-constrained devices Stelia [2025]. Consequently, improving the efficiency of these test-time compute algorithms becomes essential.
Fig. 1 illustrates how test-time scaling substantially increases computational requirements. The baseline implementation Kim et al. [2025] requires between $3\times$ and $10\times$ more compute time over regular image generation to achieve performance gains of $8\%$ to $12\%$ respectively on GenEval Ghosh et al. [2023]. In contrast, our proposed method, Verifier-Threshold, reaches the same $12\%$ improvement using just $(1/4)^{\text{th}}$ of the computational cost of the baseline, clearly demonstrating its efficiency. Furthermore, our algorithm continues to benefit from scaling, achieving gains of up to $15\%$.
The key contributions of this work are twofold: identifying the “compute-dumping” issue in the existing baseline algorithm and introducing the “Verifier-Threshold” algorithm as a solution. We validate our approach on state-of-the-art image generation models and datasets. We organize the rest of the paper as follows: Sec. 2 discusses the baseline algorithm and experimental setup. Sec. 3 shows the compute dumping phenomenon and possible solutions, including our proposed Verifier-Threshold approach. Lastly, Sec. 4 presents results, including some example images generated by the different approaches.
2 Preliminaries
2.1 Test-time Scaling and FlowRBF
Diffusion models Dhariwal and Nichol [2021], Ho et al. [2020] and flow models Labs et al. [2025] along with autoregressive models Tian et al. [2024] have become the mainstream methods for image generation. Test-time scaling for image generation basically aims to use the compute available at test-time to improve performance, either through increasing the denoising steps Salimans and Ho [2022] or searching for better noises Ma et al. [2025], Kim et al. [2025]. Noise-search in particular has shown to scale much better than denoising steps Ma et al. [2025]; hence we focus on that here. FlowRBF Kim et al. [2025] proposes a key idea of non-uniform inference-compute budget allocation across denoising steps which is a crucial advance in noise-search approaches. Using this idea, it beats all previous baseline methods such as Best-of-$N$ (which performs as good as zero-order and Search-over-Paths) from Ma et al. [2025]. Hence we use FlowRBF Kim et al. [2025] as our main baseline, and show improvements on the same flow models as proof-of-concept. However, FlowRBF’s approach of non-uniform allocation and our proposed algorithm can be easily extended to other diffusion models. Prior works are discussed in detail in Appendix B.
2.2 Experimental Setup
We use the FLUX-Schnell model Labs et al. [2025] for all our experiments similar to Kim et al. [2025]. We use ImageReward Xu et al. [2023] and VQAScore Lin et al. [2024] as our verifier reward models to guide the noise-search process. For evaluating the performance of different algorithms, we use the GenEval benchmark Ghosh et al. [2023], which consists of prompts that test for attributes such as color, number, position, etc. and helps determine an objective answer. We calculate the average percentage of correct images across different attributes as our performance metric. For the efficiency metric, we use the average wall clock duration (relative to base SDv2.1) taken for each prompt on a single AMD MI300X GPU. We also use the number of function evaluations (NFEs) to indicate the efficiency of the approach (lower is better). More details are discussed in Appendix A.
3 Verifier Threshold
3.1 Compute Dumping and Manual Budgeting
Figure 2: (a), (b) and (c) show how FlowRBF distributes NFEs (or compute budget) across denoising steps, averaged across all GenEval prompts with the total NFE budget set to 40, 80, and 160, respectively. Notice that in all cases, the inference-compute gets "dumped" at the last denoising step, and this issue persists even at total NFE budget as large as 500. (d) shows the verifier score as a function of denoising steps averaged over all GenEval prompts for total NFEs = 80. Notice that the verifier score does not increase at the last step in proportion to the amount of compute dedicated to it, thus wasting the test-time compute budget.
The efficacy of the FlowRBF Kim et al. [2025] approach can be mainly attributed to the idea of non-uniform inference-compute allocation across denoising steps or rollover budget forcing (RBF). The basic concept behind the RBF algorithm is that the denoiser stays in a particular denoising step until either: (a) the verifier score for the new noise particle exceeds the previous step’s highest score, or (b) the per-step maximum compute budget is exhausted. The issue with these criteria is that: (a) it is greedy and jumps to the next step even with a marginally better particle, and (b) it does not account for the “wrong” trajectories the diffusion process takes due to the per-step budget.
Manifestations of these issues are shown in Figs. 2 and 3 where the model performs a ten-step denoising process. Fig. 2 (a, b, c) show that even though the RBF algorithm is designed to allow non-uniform allocation of inference-compute (NFEs), most of the steps (1 to 8) are allocated similar NFEs (except for the zeroth step which is hard-coded to use a pre-defined amount of NFEs). Significant portions of the total test-time NFEs are “dumped” at the last step. When the NFE budget is lower (=40) (Fig. 2 (a)), the algorithm quickly exhausts the per-step budget and moves to the next step. When the NFE budget is higher (=80, 160) (Figs. 2 (b & c)), the greediness of the RBF algorithm causes it to quickly move to the next step. In both cases, the algorithm has a large NFE budget remaining which it is forced to "dump" in the final denoising step. Fig. 2 (d) shows that the average verifier score at the last denoising step does not increase in proportion to the allocated compute, thus underscoring the dumping issue. Fig. 3 (a) shows an example prompt that highlights the mistake caused by the compute dumping phenomenon.
Intuitively, the RBF budget allocations in Fig. 2 are sub-optimal since the structural changes necessary to obtain a semantically correct image happens in the initial denoising steps, implying that more test-time budget needs to be allocated there rather than at the end. Manually hard-coding the compute distribution to have more budget at the initial steps is one approach to solving this problem as shown in Fig. 3 (b). Although this approach may yield improvements for certain prompts, it does not generalize well across all prompts. Therefore, an algorithm that automatically redistributes the compute budget is necessary.
(e) (a)
(f) (b)
Figure 3: (a) shows three images generated at different denoising steps for the prompt “a photo of a bench left of a bear” using FlowRBF Kim et al. [2025], with a total of 40 NFEs and ImageReward Xu et al. [2023] as the verifier. This example illustrates the compute-dumping issue. The first image (at the second denoising step) captures the correct structure, though the texture remains noisy. However, the second image, produced at the third step, achieves the highest verifier score and is therefore selected, despite placing the bear on top of the bench rather than to the left. Because no additional budget is allocated at earlier steps, the denoising process continues from this incorrect latent, ultimately producing a final image with the bear on top of the bench. (b) demonstrates how this issue could be mitigated by manually reallocating the compute budget through hard-coded NFEs at each step (i.e., manual budgeting).
3.2 Verifier-Threshold Algorithm
Rather than advancing to the next denoising step as soon as a higher-reward noise particle is obtained, we propose waiting until the verifier score surpasses the previous one by a significant margin (i.e., a threshold). In this way, the algorithm progresses only when the score at a given step is sufficiently improved. This mechanism forms the basis of our Verifier-Threshold algorithm, as illustrated in Fig. 4a and detailed in Alg. 1.
Algorithm 1 Verifier-Threshold Algorithm
1:Input: compute budget NN, denoising timesteps $T$, generator model $\mathcal{G}_{\theta}$, verifier model $\mathcal{R}$, prompt $p$, Verifier-Threshold $\Delta$
2:Output: image
3:$t=1$
4:for n=1,Nn={1,...N} do
5:# sample a noise particle and perform 1 NFE
6:  $\epsilon_{n}\sim\mathcal{N}(\mathbf{0,I})$
7:  $z_{t}=\mathcal{G}_{\theta}(z_{t-1},\epsilon_{n},t,p)$
8:# calculate reward score using verifier
9:  $r_{t}=\mathcal{R}(\text{Tweedie's}(z_{t}),p)$
10:# Verifier-Threshold criteria
11:  if $r_{t}-r_{t-1}>\Delta$ : then
12:   $t++$
13:  end if
14:  if $t>T$ : then
15:   break
16:  end if
17:end for
18:return Tweedie’s ($z_{t}$)
Figure 4: (a) verifier threshold mechanism and (b) its compute allocation.
4 Results
As described in Sec. 2, we evaluate the performance of different algorithms on the GenEval benchmark Ghosh et al. [2023]. Fig. 1 highlights the advantages of our approach when using VQAScore Lin et al. [2024] as the verifier. To demonstrate that these benefits generalize across verifiers, we also evaluate performance using ImageReward Xu et al. [2023], as shown in Fig. 5. In both cases, our method achieves $2\times$ to $4\times$ higher efficiency for the same benchmark score, or alternatively, a $4\%$ to $5\%$ improvement in benchmark score at the same computational cost. All reported results correspond to test-time compute within a practical range of 2–3 minutes. Additionally, Fig. 6 presents example prompts where our algorithm produces objectively correct outputs compared to baseline methods. More comprehensive results and further examples are provided in Appendix C.
Figure 5: Test-time scaling when ImageReward Xu et al. [2023] is used as verifier.
Figure 6: The outputs of the model for the prompt “a photo of a wine glass right of a hot dog”. (a), (b), (c) represent the images generated by FLUX-Schnell Labs et al. [2025] with no test-time scaling, FlowRBF Kim et al. [2025], and Verifier-Threshold, respectively. Figures (a) & (b) have the wine glass on the left while (c) accurately provides what the prompt is requesting.
5 Conclusion
In this work, we identify a key flaw (compute dumping) in the state-of-the-art test-time scaling algorithm Kim et al. [2025] and propose the Verifier-Threshold algorithm as an effective solution. We demonstrate its performance and efficiency advantages across multiple verifiers on the GenEval benchmark. For future work, we plan to investigate adaptive verifier thresholds that vary across denoising steps and explore strategies to automate their selection. We also aim to extend our approach to standard diffusion models (beyond flow-based models).
Appendix A Experimental Setup
A.1 Dataset & Models
For evaluating our models we employ the widely-used GenEval dataset Ghosh et al. [2023]. It contains 553 prompts which checks six different attributes: single object, two object, counting, colors, position, and attribute binding. For the generator, we use the FLUX-Schnell model Labs et al. [2025] extensively due to its SOTA performance and baseline implementation from FlowRBF Kim et al. [2025]. We run it consistently for ten denoising steps for best results. For the verifiers, we use ImageReward Xu et al. [2023] and VQAScore Lin et al. [2024] which are both variants of the CLIP model Radford et al. [2021].
A.2 Methods
We use the same setup and hyperparameters as FlowRBF Kim et al. [2025] for all the methods: Best-of-$N$, search-over-paths, SVDD, and FlowRBF itself. We use a verifier threshold value ($\Delta$) of 0.005 and 0.00125 for ImageReward and VQAScore respectively when the total NFEs is 40. For subsequent total NFEs of 80 and 160, we correspondingly multiplied the threshold values with $2\times,4\times$ the above value.
Appendix B Related Work
Test-time compute as a new scaling axis
In recent years, scaling of AI models has mainly involved increasing model size and the amount of compute and data utilized during training. However, recent work shows that spending more compute at test-time can substantially improve output quality and reasoning capability, thus creating a new dimension for scaling compute to improve model performance. OpenAI’s o1 series popularized this paradigm by spending more test-time budget on deliberate “thinking” before answering OpenAI [2024]. DeepSeek-R1/R1-Zero similarly leverages extended reasoning trajectories at inference Guo et al. [2025]. Snell et al. formalize when scaling test-time compute can be more effective than scaling model size Snell et al. [2024]. Complementing these analyses, s1 demonstrates a simple and effective test-time scaling recipe with strong empirical gains Muennighoff et al. [2025].
Test-time scaling for image generation
While diffusion models and flow models have achieved impressive results in image and video generation, naive generation with these models often fails to satisfy complex user instructions involving object quantities, relative position, and size, among other conceptual attributes. Recently, test-time scaling methods have been developed for diffusion and flow models that significantly outperform naive generation.
Reward-based sampling and noise-search
Reward-based sampling methods repeatedly sample from the model’s learned distribution, guided by a reward such as text–image alignment or aesthetic quality, to select higher-quality outputs. The most basic example is Best-of-$N$ (BoN), which generates $N$ independent images with different random seeds and picks the one with the highest reward. Ma et al. (2025) formalize this paradigm by introducing both BoN and Search-over-Paths (SoP), the latter refining trajectories by sampling particles along the denoising path Ma et al. [2025]. Building on this idea, “noise trajectory search” methods incorporate reward-guided selection directly into the denoising process: SVDD Li et al. [2024a] selects, at every denoising step, the particle with the highest expected reward. Extending reward-based search to flow models, Kim et al. (2025) Kim et al. [2025] propose three complementary techniques: test-time SDE conversion, which introduces stochasticity into otherwise deterministic flows; interpolant conversion, which alters the trajectory interpolant to expand the search space; and Rollover Budget Forcing (RBF), which dynamically reallocates compute across timesteps by advancing any particle that exceeds the previous best reward and rolling over unused function evaluations to future steps, thereby ensuring efficient utilization of the total budget.
Appendix C Comprehensive Results
In this section, we flush out the plots in Fig. 1 & Fig. 5 in the form of Tab. 1. We also show the images generated for various example prompts, especially in cases where our approach succeeds when other baselines fail, and also some cases where all approaches fail. In all the example images, Fig. (a) corresponds to the image generated without test-time scaling, Fig. (b) corresponds to the one generated using FlowRBF Kim et al. [2025] and Fig. (c) is the one generated by our Verifier-Threshold algorithm.
Method Generator Verifier NFEs time ($\times$) GenEval (%)
Regular-T2I SDv2.1 - 1 1 47.93
Regular-T2I FLUX-Schnell - 10 4.74 62.33
Steering Singhal et al. [2025] SDv2.1 ImageReward Xu et al. [2023] 4 3.57 57.93
Steering Singhal et al. [2025] SDv2.1 ImageReward Xu et al. [2023] 20 15.62 64.91
Steering Singhal et al. [2025] SDv2.1 ImageReward Xu et al. [2023] 40 31.91 66.18
BoN Ma et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 40 9.70 70.62
BoN Ma et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 80 17.16 70.36
BoN Ma et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 160 26.39 72.45
BoN Ma et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 40 10.29 70.07
BoN Ma et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 80 18.72 70.54
BoN Ma et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 160 27.07 71.91
SoP Ma et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 250 21.00 60.68
SoP Ma et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 250 22.25 61.19
SVDD Li et al. [2024b] FLUX-Schnell ImageReward Xu et al. [2023] 250 43.69 68.76
SVDD Li et al. [2024b] FLUX-Schnell VQAScore Lin et al. [2024] 250 56.74 72.89
FlowRBF Kim et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 40 12.80 70.10
FlowRBF Kim et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 80 18.11 72.45
FlowRBF Kim et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 160 34.11 74.20
FlowRBF Kim et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 40 13.83 70.32
FlowRBF Kim et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 80 27.03 71.77
FlowRBF Kim et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 160 48.98 73.45
VT (Ours) FLUX-Schnell ImageReward Xu et al. [2023] 40 10.31 71.61
VT (Ours) FLUX-Schnell ImageReward Xu et al. [2023] 80 18.68 76.77
VT (Ours) FLUX-Schnell ImageReward Xu et al. [2023] 160 39.36 78.62
VT (Ours) FLUX-Schnell VQAScore Lin et al. [2024] 40 12.80 73.83
VT (Ours) FLUX-Schnell VQAScore Lin et al. [2024] 80 25.90 76.12
VT (Ours) FLUX-Schnell VQAScore Lin et al. [2024] 160 51.97 77.20
Table 1: Comparison of different approaches across models, verifiers, NFEs, along with their efficiency (wall-clock time) & performance (GenEval) scores. Methods are Best-of-NN (BoN), Search-over-Paths (SoP), Soft Value-based Decoding in Diffusion (SVDD), and Verifier-Threshold (VT).
Figure 7: “a photo of a purple wine glass and a black apple”. For all examples in this section, Fig. (a) is the image generated without test-time scaling, Fig. (b) is generated using FlowRBF Kim et al. [2025] and Fig. (c) is generated by our Verifier-Threshold algorithm.
Figure 8: “a photo of four clocks”
Figure 9: “a photo of a bench left of a bear”
Figure 10: “a photo of three sports balls”
Figure 11: “a photo of a train above a potted plant”
Figure 12: “a photo of a brown oven and a purple train”
Figure 13: “a photo of a purple carrot”
Figure 14: “a photo of a wine glass right of a hot dog”
Figure 15: “a photo of a purple computer keyboard and a red chair”
Figure 16: “a photo of a white banana and a black elephant”
Figure 17: “a photo of an orange handbag and a green carrot”
Figure 18: “a photo of an orange donut and a yellow stop sign”
Figure 19: “a photo of a sports ball left of an umbrella”